An ambulance travels 315 miles in 4 hours 30 minutes. At what average speed does it travel ?
Found 2 solutions by ewatrrr, lenny460:
Answer by ewatrrr(24785)
Answer by lenny460(1073)

An ambulance travels 315 miles in 4 hours 30 minutes. At what average speed does it travel ?
4 hours 30 minutes = 4.5 hours = 4 1/2 hours

4 1/2 = 9/2

315 / (9/2) = 315 * 2/9

315 * 2 = 630

630 / 9

Average speed:

70 mph